1)what mass of carbon dioxide will contain same number of atoms as are contained in 20 g of oxygen?


2)how much sulphur is present in an organic compound ,if 0.53 g compound gave 1.158 g of BaSO4 on analysis?

Grade:12th Pass

3 Answers

yash watwani
11 years ago

1]   20/16*6.023*10^23=7.52*10^23atoms

so7.52*10^23/6.023*10^23=1.024*10^23*44=44.924*10^23

2]

Neha Gaur
11 years ago

1)32g of O2= 2xNo atoms  (No=avogadro no.)

    20g "    "  =(2/32)x20 No=1.25 No atoms

    44g of CO2=3xNo

     1.25 No=(44/3)x1.25=18.33g of carbon dioxide

 2)molar mass of BaSO4=233g

    no. of moles of S in compound=no. of moles of S in BaSO4=NO. OF MOLES OF BaSO4 =1.158/233

    amt. of sulphur in compound=(1.158/233)x16=0.07955

 

I hope this will help you

Anjali kumari
5 years ago
As in (1) the number of moles of oxygen in 20g=5/4 moles.So, the required mass of CO2 in 5/4moles=55g.